if a planet existed whose mass twice that of earth and whose radius was 2 times greater, how much a 1.0 kg mass will weigh on its surface
a)4.4N
b)6.7N
c)2.2 N
d)13.3N
what is the solution???? 

Grade:9

1 Answers

Arun
6 years ago

For earth,

Mass of earth = M

Radius of earth = R

Mass of the body = m = 1 kg

Weight on earth is, We = GMm/R2 = mg

For the other planet,

Mass = 2M

Radius = 2R

So, weight is, Wp = G(2M)(m)/(2R)2

=> Wp = (2/4)(GMm/R2)

=> Wp = (2/4)(mg)

=> Wp = (2/4)(1 × 9.8)

=> Wp = 4.9 N

Mass of the body remains same (1 kg), the weight of the body in the planet is 4.9 N.

 

Regards

Arun (askIITians forum expert)